# Memory-Efficient 2D/3D Shape Assembly of Robot Swarms

### v1

@@ III-C Physical Experiment Results
In the physical experiment, a 3D triangle and a 3D arrow are assembled by 6 and 7 Crazyflie UAVs, respectively, within a $4\,\text{m}\times 5\,\text{m}\times 3\,\text{m}$ space, as shown in Fig. 5. The experimental parameters are consistent with those used in the 3D simulations, except for $r_{\text{sense}}=1.2\,\text{m}$, $v_{\text{max}}=0.3\,\text{m/s}$, and $d_{\text{max}}=5$. The system configuration is illustrated in Fig. 5(a), consisting of UAVs, a PC running an open-source aerial swarm control toolbox (CrazyChoir) [29], radio dongles connected to the PC, and a motion capture system. Considering the instability of peer-to-peer communication among robots, the PC is added to receive real-time robot state information from the motion capture system and execute the proposed algorithm in a distributed manner to compute the control commands. The control commands are then transmitted to the UAVs via the radio dongles at a control frequency of $20\,\text{Hz}$. Such an experimental setup is commonly used for validating distributed algorithms in physical multi-robot systems [21, 30].
During the experiments, all UAVs first take off and hover at approximately $1\,\text{m}$, then the control algorithm is activated and executed for $40\,\text{s}$. Note that the times reported in this figure refer only to the control phase, excluding take-off and hovering. Fig.5(b) shows the recorded trajectories of all robots and the entering rate curves for the two shapes. In both experiments, all robots enter the shape after approximately $20\,\text{s}$. The corresponding snapshots from the top and oblique side views are presented in Fig.5(c) and (d). The snapshots show that both shapes are assembled within $30\,\text{s}$ and maintained until $40\,\text{s}$. The discrepancy between the experimental and simulation assembly efficiency arises mainly from differences in maximum velocity constraints and control frequency due to hardware limitations, as well as real-world system dynamics such as downwash and ground effects. These results demonstrate that the UAV swarm can assemble different target shapes efficiently and robustly, thereby validating the real-world applicability of the proposed algorithm.
@@ IV Conclusion
This letter has presented a novel framework for robot swarm shape assembly in both 2D and 3D scenarios. The proposed framework integrated a memory-efficient, tree-based shape encoding-decoding scheme with a behavior-based, distributed assembly controller. Compared with a state-of-the-art image-based baseline, our approach reduced memory usage by one to two orders of magnitude, which is critical for fully distributed systems and resource-constrained robots. Furthermore, 2D and 3D simulations with up to 400 robots demonstrated dimension-independent performance, achieving two to three times faster shape entry while maintaining comparable final uniformity relative to the baseline. Finally, experiments with a swarm of 6 to 7 UAVs validated the practical feasibility of our framework. Future research will focus on eliminating the global frame assumption, reducing dependency on the motion capture system, and developing robust communication strategies to cope with intermittent links in 3D deployments.

### v2

@@ III-C Physical Experiment Results
In the physical experiment, a 3D triangle and a 3D arrow are assembled by 6 and 7 Crazyflie UAVs, respectively, within a $4\,\text{m}\times 5\,\text{m}\times 3\,\text{m}$ space, as shown in Fig. 5. The experimental parameters are consistent with those used in the 3D simulations, except for $r_{\text{sense}}=1.2\,\text{m}$, $v_{\text{max}}=0.3\,\text{m/s}$, and $d_{\text{max}}=5$. The system configuration is illustrated in Fig. 5(a), consisting of UAVs, a PC running an open-source aerial swarm control toolbox (CrazyChoir) [15], and a motion capture system. The motion capture system and the ground station PC are used as an engineering implementation to ensure reliable state acquisition and to address onboard computation constraints. They do not alter the distributed information flow: the PC runs independent per-robot control threads in parallel; each thread uses only that UAV’s current state and neighborhood information and transmits the control command to that UAV at a control frequency of $20\,\text{Hz}$. This implementation practice is consistent with the baseline method [25].
During the experiments, all UAVs first take off and hover at approximately $1\,\text{m}$, then the control algorithm is activated and executed for $40\,\text{s}$. Note that the times reported in Fig. 5 refer only to the control phase, excluding take-off and hovering. Fig. 5(b) shows the recorded trajectories of all robots and the entering rate for the two shapes. The entering rate denotes the percentage of robots inside the shape. In both experiments, all robots enter the shape after approximately $20\,\text{s}$. The corresponding snapshots from the top and oblique side views are presented in Fig. 5(c) and (d). The snapshots show that both shapes are assembled within $30\,\text{s}$ and maintained until $40\,\text{s}$. The discrepancy between the experimental and simulation assembly efficiency arises mainly from differences in maximum velocity constraints and control frequency due to hardware limitations, as well as real-world system dynamics such as downwash and ground effects. These results demonstrate that the UAV swarm can assemble different target shapes efficiently and robustly, thereby validating the real-world applicability of the proposed algorithm. Note that these experiments are intended as a feasibility validation under practical real-world constraints, whereas the primary evidence for scalability is provided by the large-scale simulations.
@@ IV Conclusion
This letter has presented a novel framework for robot swarm shape assembly in both 2D and 3D scenarios. The proposed framework integrated a memory-efficient, tree-based shape encoding-decoding scheme with a behavior-based, distributed assembly controller. Compared with a representative image-based baseline, our approach reduced memory usage by one to two orders of magnitude, which is critical for fully distributed systems and resource-constrained robots. Furthermore, 2D and 3D simulations with up to 400 robots demonstrated dimension-independent performance of our method, achieving two to four times faster shape entry compared with the baseline. Finally, experiments with a swarm of 6 to 7 UAVs validated the real-world feasibility of our framework. Future research will focus on eliminating the global frame assumption, mitigating the downwash effect, reducing dependency on the motion capture system, and developing robust communication strategies to cope with intermittent links in 3D deployments.